OpenStudy (anonymous):

Use addition or substitution to find the value of x for this set of equations. 2x - 3y = -15 x = 4y

OpenStudy (hba):

Multiply eq(2) by 2 and subract them :)

OpenStudy (hba):

@Corinne202 Got it ?

OpenStudy (anonymous):

Do I multiply both equations by 2?

OpenStudy (ajprincess):

nope only the second.:)

OpenStudy (hba):

Only the second :)

OpenStudy (anonymous):

How do I find out what x is doing that

OpenStudy (hba):

2x - 3y = -15---(1) x = 4y------(2) 1)Multiply (2) by 2 2)Subract (1) with (2)

OpenStudy (hba):

@Corinne202 Is that understood ?

OpenStudy (anonymous):

Yeah :)

OpenStudy (hba):

Good to hear :)

OpenStudy (hba):

What did you get ?

OpenStudy (anonymous):

When I subtracted them I got y=3 then I put that in the first equation to find x= -12, is that correct?

OpenStudy (hba):

y will be -3 ?

OpenStudy (hba):

2x - 3y = -15---(1) 2x-8y = 0------(2) - + ---------------- 5y=-15 y=-15/5 y=-3 Now put this value in one of the equations and find x :)

OpenStudy (hba):

@Corinne202 Understood ?